OpenStudy (firejay5):

Find the value of each variable. If your answer is NOT an integer, leave it in simplest radical form. Show work!!! Medal will be rewarded for patronage! :D

OpenStudy (firejay5):



OpenStudy (cwrw238):

since the triangle is isosceles the 2 sides including the right angle are equal sin45 = x / sqrt6 x = sqrt6 * sin 45 the other side is same length

OpenStudy (firejay5):

we haven't learned sin in Geometry

OpenStudy (cwrw238):

oh ok then use the Pythagoras theorem let the other side be = x then (sqrt6)^2 = x^2 + x^2 2 x^2 = 6 can you continue?

OpenStudy (firejay5):

sqrt of 3

OpenStudy (cwrw238):

yep

OpenStudy (firejay5):

x = sqrt of 3

OpenStudy (cwrw238):

yes, the 2 sides including the right angle are both = sqrt3

OpenStudy (jdoe0001):

*cough* http://grockit.com/blog/collegeprep/files/2011/04/35.png *cough*

OpenStudy (cwrw238):

yes - its worth remember the 2 triangles in the above link

OpenStudy (cwrw238):

OK? do you need any more explanation?

OpenStudy (firejay5):

Nope so problems like this, I can use Pythagorean Theorem

OpenStudy (cwrw238):

yes - but you can also solve them quickly by committing to memory the 2 triangles that jdoe has mentioned in in his post

OpenStudy (cwrw238):

though I'm fascinated by his 'cough' - lol.